Draw all possible distinct sample of size 3 from the population 2 ,4, 6, 8, 10, 12,




Construct the sampling distribution of the sample mean X and calculate it's means and variance meauX and sigma X

1
Expert's answer
2022-01-18T18:29:26-0500

Mean


"\mu=\dfrac{2+4+6+8+10+12}{6}=7"

Variance


"\sigma^2=\dfrac{1}{6}\big((2-7)^2+(4-7)^2+(6-7)^2""+(8-7)^2+(10-7)^2+(12-7)^2\big)""=\dfrac{70}{6}=\dfrac{35}{3}"

Standard deviation


"\\sigma=\\sqrt{\\sigma^2}=\\sqrt{\\dfrac{35}{3}}\\approx\t3.415650"



We have population values "2,4,6,8,10,12" population size "N=6" and sample size "n=3." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{6}{3}=\dfrac{6!}{3!(6-3)!}=20""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,4,6 & 4 \\
 \hdashline
 2 & 2,4,8 & 14/3 \\
 \hdashline
 3 & 2,4,10 & 16/3 \\
 \hdashline
 4 & 2,4,12 & 6 \\
 \hdashline
 5 & 2,6,8 & 16/3 \\
 \hdashline
 6 & 2,6,10 & 6 \\
 \hdashline
 7 & 2,6,12 & 20/3 \\
 \hdashline
 8 & 2,8,10 & 20/3 \\
 \hdashline
 9 & 2,8,12 & 22/3 \\
 \hdashline
 10 & 2,10,12 & 8 \\ 
 \hdashline
 11 & 4,6,8 & 6 \\ 
 \hdashline
 12 & 4,6,10 & 20/3 \\ 
 \hdashline
 13 & 4,6,12 & 22/3 \\ 
 \hdashline
 14 & 4,8,10 & 22/3 \\ 
 \hdashline
 15 & 4,8,12 & 8 \\ 
 \hdashline
 16 & 4,10,12 & 26/3 \\ 
 \hdashline
 17 & 6,8,10 & 8 \\ 
 \hdashline
 18 & 6,8,12 & 26/3 \\ 
 \hdashline
 19 & 6,10,12 & 28/3 \\ 
 \hdashline
 20 & 8,10,12 & 10 \\ 
 \hdashline
\end{array}"



The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 4 & 1& 1/20 & 12/60 & 144/180 \\
 \hdashline
 14/3 & 1 & 1/20 & 14/60 & 196/180 \\
 \hdashline
 16/3 & 2 & 2/20 & 32/60 & 512/180 \\
 \hdashline
 6 & 3 & 3/20 & 54/60 & 972/180\\
 \hdashline
 20/3 & 3 & 3/20 & 60/60 & 1200/180 \\
 \hdashline
 22/3 & 3 & 3/20 & 66/60 & 1452/180 \\
 \hdashline
 8 & 3 & 3/20 & 72/60 & 1728/180 \\
 \hdashline
 26/3 & 2 & 2/20 & 52/60 & 1352/180 \\
 \hdashline
 28/3 & 1 & 1/20 & 28/60 & 784/180 \\
 \hdashline
 10 & 1 & 1/20 & 30/60 & 900/180 \\
 \hdashline
 Total & 20 & 1 & 420/60 & 9240/180 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{420}{60}=7"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=\mu_{\bar{x}}=7=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2""=\dfrac{9240}{180}-(7)^2=\dfrac{7}{3}""\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{\dfrac{7}{3}}\approx1.527525"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{35}{3(3)}(\dfrac{6-3}{6-1})""=\dfrac{7}{3}, True"